a student performs an experiment to find average acceleration of a falling object. He drops a baseball from a building and uses a string and meter stick to measure the height the ball was dropped from. He uses a stopwatch to find an average time of fall for 3 trials from the same height and reports the following data: h=5.25+or- 0.15m t=1.14+or-0.06s . Use the equation a=2h/t to determine the average acceleration and its uncertainity
1
Expert's answer
2018-09-24T12:25:08-0400
The average acceleration:
a=2h/t^2 =2(5.25)/(1.14)^2 =8.1 m/s^2 .
Δh/h=0.15/5.25=0.029; 2 Δt/t=2 0.06/1.14=0.106.
Thus,
Δa/a=√(〖0.029〗^2+〖0.106〗^2 )=0.109.
Δa=0.109(8.1)=0.9 m/s^2 .